OpenStudy (anonymous):

If a 75-kg skater starts his skate at 8.0m, at his lowest point (height = 0), what is his velocity?

OpenStudy (anonymous):

mgh= 1/2 m v^2 gh= 1/2 v^2 9.8 * 8= 1/2 v^2 solve for v

OpenStudy (lalaly):

V = sqrt(2gH) where H = 8 m.

OpenStudy (anonymous):

thanks for the help! Awesome!

OpenStudy (anonymous):

If a 75-kg skater starts his skate at 8.0m, at his lowest point (height = 0), what is his velocity?